# NSPDI-SNN: An efficient lightweight SNN based on nonlinear synaptic pruning and dendritic integration

### v1

@@ Abstract
[label4]organization=Research Unit of NeuroInformation (2019RU035), city=Chengdu, postcode=621731, state=Sichuan, country=China
@@ 1 Introduction
However, simply copying the development model of ANNs does not take full advantage of SNNs. Using the large model as ANNs also makes SNN encounter the problem of the computing power bottleneck. Therefore, further research needs to be considered from more dimensions. Biological neural networks are more similar to SNNs than ANNs and offer superior reliability and efficiency. Simulating the properties of biological neurons has the potential to enhance the performance of SNNs. A considerable number of brain-inspired studies have been published. Among these, a particularly effective method inspired by neuronal heterogeneity involves parameterizing or dynamizing the membrane potential constants (Fang et al., 2021) or firing thresholds (Sun et al., 2023; Ding et al., 2022) of leaky integrate-and-fire (LIF) neurons (Dayan et al., 2002). Researchers have shown that, unlike the dendrites of simple point neurons, those of pyramidal neurons enable more complex functions, such as ”XOR” arithmetic (Gidon et al., 2020). In additional studies, researchers have shown that the activity of individual pyramidal neurons is comparable to that of deep neural networks (Beniaguev et al., 2021). Therefore, a practical approach to improve deep neural networks may be to model the complex structure and functions of dendrites.
@@
A recent biologically inspired pruning technique, known as the state transition of dendritic spines (STDS), simulates synaptic sparsification by by reparameterizing connection weights (Chen et al., 2022). STDS is inspired by the state transitions between mature dendritic spines-filopodia and excitatory-inhibitory synapses observed in biological neurons, especially in Purkinje cells. In our approach, we further extend this concept by explicitly modeling the heterogeneity of state transition ratios ( (Dobrunz and Stevens, 1997; Fritschy et al., 2012; Perez-Nieves et al., 2021)) in different neurons and neuronal populations. This neurobiological variability is a foundation for developing a more flexible and adaptive pruning mechanism, as illustrated in Fig. 1(a). To further enhance sparsity, we incorporate a threshold-based pruning mechanism with the state transition ratio. This dual-pruning strategy enables the synaptic reparameterization process to adapt more flexibly to structural heterogeneity, allowing the network to learn diverse and effective patterns of synaptic transformation grounded in biological plausibility.
@@ 2.1 Nonlinear Dendritic Integration
Dendrites provide neurons with powerful computational capabilities. Learning from dendritic structures is an approach to improve AI performance. One method is polynomial neural networks (Ma and Khorasani, 2005; Jiang et al., 2016), which integrates high-order function formulas into neuron activation functions, thereby improving the nonlinearity of neurons. Another method directly introduces higher-order term relationships between input information to replace the traditional neuron structure (Chen and Liu, 2022; Liu and Wang, 2022). It has been verified on many simple datasets, but further exploration is required on complex datasets. A common method is incorporating nonlinear calculations into the input integration process (Mantini and Shah, 2021; Xu et al., 2022). Nonlinear calculations greatly increase the computational complexity and parameter amount of the model and can improve performance on traditional datasets. In recent experimental calculations and observations, the researchers have found that dendritic computations that involve spiking adhere to a bilinear process (Hao et al., 2009; Li et al., 2014, 2019). Based on this discovery, we propose integrating the bilinear property into SNN computations to enhance computational performance when a highly sparse model is preserved.
@@ 3.2.1 State Transition of Heterogeneous Dendritic Spines
$\bm{W}_{\mathrm{p}}=\operatorname{sign}\left(\bm{\Theta}\right)\odot\left[\bm{A}\odot\left(|\bm{\Theta}|-d_{1}\right)\right]_{+},d_{1}\geq 0,$ (7)
where $\bm{W}_{\mathrm{p}}\in\mathbb{R}^{H\times W}$ is the post-transition weight with a size of $H\times W$, $\bm{A}\in\mathbb{R}^{H\times W}$ is the learnable transition gain matrix, which represents the extent to which the parameter $\bm{\Theta}\in\mathbb{R}^{H\times W}$ effects the real connection weight $\bm{W}_{\mathrm{p}}$; and $d_{1}\in\mathbb{R}$ signifies the transition threshold. $\bm{A}$ can be hierarchically defined at three levels, analogous to nonlinear integration: synapse-level, channel-level (neuron-level for fully connected layers), and layer-level. Its mathematical formulation follows a similar structure to the nonlinear integration weight $\bm{V}$, and thus, for brevity, we omit the detailed derivation here.
Connections with values exceeding threshold $d_{1}$ (i.e., $\{\theta\in\bm{\Theta}\mid\theta>d_{1}\}$) are classified as dendritic spine connections, while the remaining connections ($\{\theta\in\bm{\Theta}\mid\theta\leq d_{1}\}$) correspond to filopodia. Increasing $d_{1}$ reduces the proportion of mature spines and decreases the connection weights, thus the network becomes sparser. Conversely, when $d_{1}$ decreases, the proportion of dendritic spines increases and the connection weights also increase, thus the network becomes denser. The strength of synaptic connections varies because of differences in $\bm{A}$. This natural variation lets the remaining connections (after pruning) make up for lost information by adjusting their strengths, similar to how synaptic scaling works in real neurons to keep networks stable.
@@ 3.3 Learning Method
$\nabla_{w}\mathcal{L}$ $=\nabla_{I}\mathcal{L}\cdot\frac{\partial I}{\partial w}=\nabla_{I}\mathcal{L}\cdot s,$ (13)
$\nabla_{v}\mathcal{L}$ $=\nabla_{I}\mathcal{L}\cdot\frac{\partial I}{\partial v}=\sigma\cdot\nabla_{I}\mathcal{L}+\nabla_{I}\mathcal{L}\cdot s,$
@@
Dataset Trial Epoch BS $\eta$ $\gamma_{\eta}$ $u_{\mathrm{t}h}$ $\tau_{m}$ $T$ $\Delta t(\mathrm{ms})$
DVS128 Gesture 10 500 32 1e-3 Cosine 1.0 3.3 10 125
CIFAR10-DVS 10 200 80 2e-4 Cosine 1.0 2.0 20 50
CIFAR10 3 400 100 5e-2 Cosine 1.0 2.0 4 -
SHD 10 200 128 1e-2 Cosine 0.3 2.0 100 14
Maze2D 10 - 100 1e-3 - 0.5 2.0 4 -
@@
Based on Eqs. (7) and (8), gradient descent can directly optimize $\theta$ and $a$. The gradient of the reparameterization parameter $\nabla_{\theta}\mathcal{L}$ and the gradient of the transition gain $\nabla_{a}\mathcal{L}$ are zero when $\left|\theta\right|\leq\frac{d_{1}+d_{2}}{a}$; in all other cases, they can be described as follows:
and
$\nabla_{a}\mathcal{L}$ $=\nabla_{v}\mathcal{L}\cdot\frac{\partial v}{\partial a}=\nabla_{v}\mathcal{L}\cdot\theta,$ (15)
$\nabla_{\theta}\mathcal{L}$ $=\nabla_{v}\mathcal{L}\cdot\frac{\partial v}{\partial\theta}=\nabla_{v}\mathcal{L}\cdot a.$
Using Eqs. (10)-(15), NSPDI-SNN enables efficient pruning during training. When the NSP or NDI method is removed, our model degenerates into a dense SNN based on NDI (NDI-SNN) or lightweight SNN based on NSP (NSP-SNN), respectively.
@@ 4.1 Datasets and Implementations
All experiments were implemented using the PyTorch framework and executed on servers equipped with NVIDIA A100 GPUs. Unless otherwise noted, we adopted the default hyperparameter configuration outlined in Table 1, including the number of trials, epoch, batch size (BS), learning rate ($\eta$), learning rate decay schedule ($\gamma_{\eta}$), membrane time constant ($\tau_{\mathbf{m}}$), neuron firing threshold ($u_{\mathrm{th}}$), simulation time window ($T$), and integration timestep ($\Delta t$). All reported performance metrics represent the mean and standard deviation of the top-1 classification accuracy across multiple independent runs, along with the corresponding best-performing scores.
@@
Importantly, as the classification head plays a critical role in determining final performance, we excluded it from NSPDI pruning to preserve its full representational capacity. The pruning schedule for $d_{1}$ followed the sinusoidal scheme introduced in STDS, while $d_{2}$ was governed by a multi-periodic sine schedule inspired by cyclical pruning strategies (Srinivas et al., 2022).
@@
Figure 4: Accuracy difference with respect to the baseline sparse model (Base + STDS) under varying sparsity levels on three datasets. Each line corresponds to one model: NSPDI (red), Base + NSP(yellow), and NDI + STDS (green). (a) DVS128 Gesture dataset. (b) CIFAR10-DVS dataset. (c) CIFAR10 dataset.
@@ 4.2 NDI Improved Spatiotemporal Representation in SNNs
In contrast, for the static CIFAR10 dataset, where temporal information is absent, the performance gains are comparatively modest. The average accuracy remains around 96%, indicating that NDI’s advantage is less pronounced in temporally invariant settings. These findings collectively confirm the utility of NDI, particularly in domains rich in spatiotemporal complexity.
@@
Figure 5: Accuracy comparison of NSPDI-SNN models using transition gains introduced at different structural levels: synapse-level (S, blue), channel-level(C, teal), and layer-level (L, green) levels. Each plot corresponds to one dataset: (a) DVS128 Gesture. (b) CIFAR10-DVS. (c) CIFAR10.
@@ 4.3 Ablation Study with NSP
To enhance computational efficiency and reduce parameter redundancy in the high-performance NDI-SNN architecture, we introduce a lightweight pruning strategy termed NSP. NSP is an optimized extension of the STDS pruning framework, tailored to improve sparsity-aware training in spiking neural networks. To systematically evaluate the contribution of NSP under sparse learning, we applied both STDS and NSP to prune two model variants: the baseline SNN and the NDI-SNN. This yielded four distinct configurations: base + STDS, NDI + STDS, base + NSP, and NSPDI. All pruning experiments were conducted with a fixed pruning schedule, where the initial pruning threshold $d_{1}$ was set to 100%, and the enhancement threshold $d_{2}$ was set to 0.
Figure 4 illustrates the accuracy improvements of base + NSP, NDI + STDS, and NSPDI, reported relative to the baseline sparse model (base + STDS) across three datasets: DVS128 Gesture, CIFAR10-DVS, and CIFAR10. NSP consistently outperforms STDS when applied to the baseline model, confirming its efficacy in enhancing sparse model performance. More notably, NSPDI achieves the highest accuracy across all evaluated sparsity levels. On the DVS128 Gesture dataset, NSPDI demonstrates a significant performance margin, particularly at higher sparsity levels, indicating its robustness in preserving spatiotemporal dynamics under structural compression. Similarly, on CIFAR10-DVS and CIFAR10, NSPDI maintains superior accuracy across nearly all sparsity regimes, underscoring the effectiveness of NSP in complementing the representational benefits introduced by NDI. These results validate the synergistic integration of NSP and NDI, highlighting NSPDI as a compact yet high-performing solution capable of balancing accuracy and computational efficiency in static and event-driven domains.

### v2

@@ Abstract
[label4]organization=Research Unit of NeuroInformation (2019RU035), city=Chengdu, postcode=611731, state=Sichuan, country=China
@@ 1 Introduction
However, simply copying the development model of ANN does not take full advantage of SNNs. Using the large model as ANNs also makes SNN encounter the problem of the computing power bottleneck. Therefore, further research needs to be considered from more dimensions. Biological neural networks are more similar to SNNs than ANNs and offer superior reliability and efficiency. Simulating the properties of biological neurons has the potential to enhance the performance of SNNs. A considerable number of brain-inspired studies have been published. Among these, a particularly effective method inspired by neuronal heterogeneity involves parameterizing or dynamizing the membrane potential constants (Fang et al., 2021) or firing thresholds (Sun et al., 2023; Ding et al., 2022) of leaky integrate-and-fire (LIF) neurons (Dayan et al., 2002). Researchers have shown that, unlike the dendrites of simple point neurons, those of pyramidal neurons enable more complex functions, such as ”XOR” arithmetic (Gidon et al., 2020). In additional studies, researchers have shown that the activity of individual pyramidal neurons is comparable to that of deep neural networks (Beniaguev et al., 2021). Therefore, a practical approach to improve deep neural networks may be to model the complex structure and functions of dendrites.
@@
A recent biologically inspired pruning technique, known as the state transition of dendritic spines (STDS), simulates synaptic sparsification by reparameterizing connection weights (Chen et al., 2022). STDS is inspired by the state transitions between mature dendritic spines-filopodia and excitatory-inhibitory synapses observed in biological neurons, especially in Purkinje cells. In our approach, we further extend this concept by explicitly modeling the heterogeneity of state transition ratios ( (Dobrunz and Stevens, 1997; Fritschy et al., 2012; Perez-Nieves et al., 2021)) in different neurons and neuronal populations. This neurobiological variability is a foundation for developing a more flexible and adaptive pruning mechanism, as illustrated in Fig. 1(a). To further enhance sparsity, we incorporate a threshold-based pruning mechanism with the state transition ratio. This dual-pruning strategy enables the synaptic reparameterization process to adapt more flexibly to structural heterogeneity, allowing the network to learn diverse and effective patterns of synaptic transformation grounded in biological plausibility.
@@ 2.1 Nonlinear Dendritic Integration
Dendrites provide neurons with powerful computational capabilities. Learning from dendritic structures is an approach to improve AI performance. One method is polynomial neural networks (Ma and Khorasani, 2005; Jiang et al., 2016), which integrates high-order function formulas into neuron activation functions, thereby improving the nonlinearity of neurons. Another method directly introduces higher-order term relationships between input information to replace the traditional neuron structure (Chen and Liu, 2022; Liu and Wang, 2022; Liu et al., 2024). However, this approach has so far only been validated on simple datasets and requires further investigation on more complex datasets. A common approach is incorporating nonlinear calculations into the input integration process (Mantini and Shah, 2021; Xu et al., 2022; Zheng et al., 2024). Nonlinear calculations greatly increase the computational complexity and parameter amount of the model and can improve performance on traditional datasets. In recent experimental calculations and observations, the researchers have found that dendritic computations that involve spiking adhere to a bilinear process (Hao et al., 2009; Li et al., 2014, 2019). Based on this discovery, we propose integrating the bilinear property into SNN computations to enhance computational performance when a highly sparse model is preserved.
@@ 3.2.1 State Transition of Heterogeneous Dendritic Spines
$w_{p}=\operatorname{sign}\left(\theta\right)\cdot\left[a\left(|\theta|-d_{1}\right)\right]_{+},d_{1}\geq 0,$ (7)
where $w_{p}\in\mathbb{R}$ is the post-transition weight. $a>0$ is the learnable transition gain. It shows how the parameter $\theta\in\mathbb{R}$ affects the real connection weight $w_{p}$. The parameter $d_{1}\in\mathbb{R}$ signifies the transition threshold. Here, $\bm{\Theta}$ denotes the set of all learnable parameters $\theta$. The $a$ can be hierarchically defined at three levels, analogous to nonlinear integration: synapse-level, channel-level (neuron-level for fully connected layers), and layer-level. Its mathematical formulation follows a similar structure to the nonlinear integration weight $\bm{V}$, and thus, for brevity, we omit the detailed derivation here.
Connections with values exceeding threshold $d_{1}$ (i.e., $\{\theta\in\bm{\Theta}\mid\theta>d_{1}\}$)) are classified as dendritic spine connections, while the remaining connections correspond to filopodia. Increasing $d_{1}$ reduces the proportion of mature spines and decreases the connection weights, thus the network becomes sparser. Conversely, when $d_{1}$ decreases, the proportion of dendritic spines increases and the connection weights also increase, thus the network becomes denser.
The strength of synaptic connections varies because of differences in $a$. In our design, $a$ is strictly positive to ensure that the post-transition weight $w_{p}$ preserves the correct sign and magnitude relative to $\theta$. During training, we enforce this constraint by clamping the updates of $a$, preventing it from becoming negative. This natural variation lets the remaining connections (after pruning) make up for lost information by adjusting their strengths, similar to how synaptic scaling works in real neurons to keep networks stable.
@@ 3.3 Learning Method
$\nabla_{w}\mathcal{L}$ $={I}\mathcal{L}\cdot\frac{\partial I}{\partial w}=\nabla_{I}\mathcal{L}\cdot x+\left(v\cdot x\right)\cdot x\cdot\nabla_{I}\mathcal{L},$ (13)
$\nabla_{v}\mathcal{L}$ $=\nabla_{I}\mathcal{L}\cdot\frac{\partial I}{\partial v}=\left(w\cdot x\right)\cdot x\cdot\nabla_{I}\mathcal{L},$
@@
Table 1: Computational complexity comparison of different modules.
Module Computing Complexity Number of parameters
Linear $O(MN)$ $O(MN)$
NDI-Linear $O(2\cdot MN+2M)$ $O(2\cdot MN+M)$
NSPDI-Linear $O(2\beta\cdot MN+2M)$ $O(2\beta\cdot MN+3\cdot M)$
Conv $O(C_{i}C_{o}K^{2}HW)$ $O(C_{i}C_{o}K^{2})$
NDI-Conv $O(2\cdot C_{i}C_{o}K^{2}HW+2\cdot C_{o}HW)$ $O(2\cdot C_{i}C_{o}K^{2}+C_{o})$
NSPDI-Conv $O(2\beta\cdot C_{i}C_{o}K^{2}HW+2\cdot C_{o}HW)$ $O(2\beta\cdot C_{i}C_{o}K^{2}+3\cdot C_{o})$
Based on Eqs. (7) and (8), gradient descent can directly optimize $\theta$ and $a$. The gradient of the reparameterization parameter $\nabla_{\theta}\mathcal{L}$ and the gradient of the transition gain $\nabla_{a}\mathcal{L}$ are zero when $\left|\theta\right|\leq d_{1}+\frac{d_{2}}{a}$. In all other cases, they can be described as follows:
and
$\nabla_{a}\mathcal{L}$ $=\nabla_{v}\mathcal{L}\cdot\frac{\partial v}{\partial a}=\nabla_{v}\mathcal{L}\cdot\theta_{v},$ (15)
$\nabla_{\theta}\mathcal{L}$ $=\nabla_{v}\mathcal{L}\cdot\frac{\partial v}{\partial\theta}=\nabla_{v}\mathcal{L}\cdot a_{v}.$
where $\theta_{w}$ and $\theta_{v}$ are the linear reparameterization weight and nonlinear reparameterization weight, respectively, and $a_{w}$ and $a_{v}$ are the linear transition gain and nonlinear transition gain, respectively.
Algorithm 1 Training procedure of NSPDI-SNN model
Input: Training dataset $\mathcal{D}_{train}$, network structure specification $\mathcal{S}$, transition threshold $d_{1}$, pruning threshold $d_{2}$
Parameter: Linear reparameterization weight $\theta_{w}$, nonlinear reparameterization weight $\theta_{v}$, transition gain $a_{w}$ and $a_{v}$, batch normalization $\mathcal{BN}$ and loss function MSELoss
Output: Trained NSPDI-SNN model
 Initialize model $\mathcal{M}$ with NSPDI-SNN layers from $\mathcal{S}$
 Load $\mathcal{D}_{train}$
for epoch do
  Update $d_{1},d_{2}$ according to epoch
  for mini-batch $(x,y)\in\mathcal{D}_{train}$ do
   for each layer $\ell$ in $\mathcal{M}$ do
    $w\leftarrow\text{NSP}(\theta_{w},a_{w},d_{1},d_{2})$,   $v\leftarrow\text{NSP}(\theta_{v},a_{v},d_{1},d_{2})$ {NSP reparameterization}. Eq. (9)
    $I\leftarrow\text{NDI}(x,w,v,b)$ {NDI integration}. Eq. (4)
    $I\leftarrow\mathcal{BN}(I)$ {Batch normal}.
    $(s,u)\leftarrow\text{LIF}(I,s,u)$ {LIF neuron update}. Eq. (1)
   end for
   $I^{o}\leftarrow\text{Linear}(s,w,b)$
   $(s^{o},u^{o})\leftarrow\text{LIF}(I^{o},s,u^{o})$ {Spiking outputs}. Eq. (1)
   $\mathcal{L}\leftarrow\text{MSELoss}(s^{o},y)$. Eq. (11)
   Backpropagate $\nabla\mathcal{L}$ and update. Eq. 10 and Eqs. (12)-(15)
  end for
end for
Using Eqs. (10)-(15), NSPDI-SNN enables efficient pruning during training. The overall training procedure is summarized in Algorithm 1. When the NSP or NDI method is removed, our model degenerates into a dense SNN based on NDI (NDI-SNN) or lightweight SNN based on NSP (NSP-SNN), respectively.
@@
Dataset Trial Epoch BS $\eta$ $\gamma_{\eta}$ $u_{\mathrm{t}h}$ $\tau_{m}$ $T$ $\Delta t(\mathrm{ms})$ $\text{Epoch}_{\text{C}}$ $T_{\text{cyc}}$
DVS128 Gesture 10 500 32 1e-3 Cosine 1.0 3.3 10 125 50 50
CIFAR10-DVS 10 200 80 2e-4 Cosine 1.0 2.0 20 50 25 25
CIFAR10 3 400 100 5e-2 Cosine 1.0 2.0 4 - 50 25
SHD 10 200 128 1e-2 Cosine 0.3 2.0 100 14 25 25
Maze2D 10 - 100 1e-3 - 0.5 2.0 4 - 25 25
@@
3.4 Computational Complexity Analysis
To systematically evaluate the computational efficiency of the proposed NSPDI method, we conduct a theoretical analysis of its computational complexity and parameter count. As shown in Tab. 1, the analysis reveals the inherent characteristics and potential advantages of the NSPDI module in terms of computational resource consumption.
The results indicate that the NSPDI module effectively reduces the computational burden for sparsity $\beta\in(0,1]$. Specifically, for the convolution module, the computational complexity of NSPDI-Conv is $O(2\beta\cdot C_{i}C_{o}K^{2}HW+2\cdot C_{o}HW)$, which reduces the dominant term of the dual-path convolution by approximately $1/\beta$ compared with NDI-Conv’s $O(2\cdot C_{i}C_{o}K^{2}HW+2\cdot C_{o}HW)$, since each path only processes $\beta\cdot C_{i}$ input channels. When $\beta<0.5$, the theoretical computational cost of NSPDI-Conv is significantly lower than that of NDI-Conv. Regarding the number of parameters, NSPDI-Conv requires $O(2\beta\cdot C_{i}C_{o}K^{2}+3\cdot C_{o})$, also approximately $1/\beta$ times smaller than NDI-Conv’s $O(2\cdot C_{i}C_{o}K^{2}+C_{o})$. NSPDI-Linear exhibits a similar optimization trend for linear modules. In summary, the theoretical analysis confirms that the NSPDI module, through a sparse connection strategy, effectively controls both computational complexity and parameter count while maintaining excellent theoretical efficiency.
@@ 4.1 Datasets and Implementations
All experiments were implemented using the PyTorch framework and executed on servers equipped with NVIDIA A100 GPUs. Unless otherwise noted, we adopted the default hyperparameter configuration outlined in Table 1, including the number of trials, epoch, batch size (BS), learning rate ($\eta$), learning rate decay schedule ($\gamma_{\eta}$), membrane time constant ($\tau_{\mathbf{m}}$), neuron firing threshold ($u_{\mathrm{th}}$), simulation time window ($T$), and integration timestep ($\Delta t$). In addition, we report two training-strategy parameters: $\text{Epoch}_{\text{C}}$ and $T_{\text{cyc}}$. Here, $\text{Epoch}_{\text{C}}$ denotes the additional epochs used after the main training phase to ensure final convergence (during these epochs $d_{1}$ and $d_{2}$ are fixed to their target values), and $T_{\text{cyc}}$ denotes the length (in epochs) of one cycle in the cyclic pruning schedule for $d_{2}$ (default $T_{\text{cyc}}=50$).
The pruning schedule for $d_{1}$ follows the sinusoidal scheme introduced in STDS. In practice, $d_{1}$ is expressed as a percentage. This directly determines the proportion of connections to prune, allowing precise and dynamic control of sparsity during training. The threshold $d_{2}$ is governed by a multi-periodic sine schedule inspired by cyclical pruning strategies (Srinivas et al., 2022). In our implementation, $d_{2}$ varies smoothly with epoch $e$ according to a (multi-)sinusoidal function with base period $T_{\text{cyc}}$ (default 50 epochs). This causes $d_{2}$ to progressively increase and decrease within each cycle. Additionally, at the end of the whole training run, we perform $\text{Epoch}_{\text{C}}$ extra epochs with $d_{1}$ and $d_{2}$ held at their final target values to guarantee convergence. This design makes pruning gradual and reversible within cycles while ensuring stability before the next cycle or at final convergence.
@@
Importantly, as the classification head plays a critical role in determining final performance, we excluded it from NSPDI pruning to preserve its full representational capacity. All reported performance metrics are the mean and standard deviation of top-1 accuracy over multiple independent runs, together with the corresponding best-performing scores.
@@
Figure 4: Accuracy difference with respect to the baseline sparse model (Linear + STDS) under varying sparsity levels on three datasets. Each line corresponds to one model: NSPDI (red), Linear + NSP(yellow), and NDI + STDS (green). (a) DVS128 Gesture dataset. (b) CIFAR10-DVS dataset. (c) CIFAR10 dataset.
@@ 4.2 NDI Improved Spatiotemporal Representation in SNNs
In contrast, for the static CIFAR10 dataset, where temporal information is absent, the performance gains are comparatively modest. The average accuracy remains around 96%, indicating that NDI’s advantage is less pronounced in temporally invariant settings. This is because the effectiveness of NDI relies on the presence of temporal dynamics, where dendritic integration can capture and exploit cross-time dependencies. In purely static datasets such as CIFAR10, these spatiotemporal dependencies degenerate into spatial correlations only, thereby limiting the potential benefits of NDI. These findings collectively confirm the utility of NDI, particularly in domains rich in spatiotemporal complexity.
@@
Figure 5: Accuracy comparison of NSPDI-SNN models using transition gains introduced at different structural levels: synapse-level (S, blue), channel-level(C, teal), and layer-level (L, green) levels. The red dashed line represents the accuracy of NDI-SNN using the STDS method. Each plot corresponds to one dataset: (a) DVS128 Gesture. (b) CIFAR10-DVS. (c) CIFAR10.
@@ 4.3 Ablation Study with NSP
To enhance computational efficiency and reduce parameter redundancy in the high-performance NDI-SNN architecture, we introduce a lightweight pruning strategy termed NSP. NSP is an optimized extension of the STDS pruning framework, tailored to improve sparsity-aware training in spiking neural networks. To systematically evaluate the contribution of NSP under sparse learning, we applied both STDS and NSP to prune two model variants: the baseline SNN and the NDI-SNN. This yielded four distinct configurations: Linear + STDS, NDI + STDS, Linear + NSP, and NSPDI. The NDI method uses the channel level, and all pruning experiments were conducted with a fixed pruning schedule, where the initial pruning threshold $d_{1}$ was set to 100%, and the enhancement threshold $d_{2}$ was set to 0.
Figure 4 illustrates the accuracy improvements of Linear + NSP, NDI + STDS, and NSPDI, reported relative to the baseline sparse model (Linear + STDS) across three datasets: DVS128 Gesture, CIFAR10-DVS, and CIFAR10. The horizontal axis represents the retention weight ratio (0.8, 0.93, 0.96, 0.99, and 0.996), and the vertical axis represents the accuracy improvement relative to the baseline sparse model (Linear + STDS). NSP consistently outperforms STDS when applied to the baseline model, confirming its efficacy in enhancing sparse model performance. More notably, NSPDI achieves the highest accuracy across all evaluated sparsity levels. On the DVS128 Gesture dataset, NSPDI demonstrates a significant performance margin, particularly at higher sparsity levels, indicating its robustness in preserving spatiotemporal dynamics under structural compression. Similarly, on CIFAR10-DVS and CIFAR10, NSPDI maintains superior accuracy across nearly all sparsity regimes, underscoring the effectiveness of NSP in complementing the representational benefits introduced by NDI. These results validate the synergistic integration of NSP and NDI, highlighting NSPDI as a compact yet high-performing solution capable of balancing accuracy and computational efficiency in static and event-driven domains.